OpenStudy (anonymous):

Integral of sqrt(64-x^2)?

OpenStudy (anonymous):

\[\int\limits \sqrt{64-x ^{2}}dx\]

OpenStudy (anonymous):

Try the trig substitution \(x=8\sin\theta\). Do you think you can take things from here? :-)

OpenStudy (anonymous):

, no why would that help?

OpenStudy (anonymous):

oh doh, what @ChristopherToni said

OpenStudy (anonymous):

I'm not very good with trig identities :/

OpenStudy (anonymous):

try it and see, it is the trick to get rid of the square root probably saw this back in trig class and figured you would never use it

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (anonymous):

Let's see why this would be case: If \(x=8\sin \theta\), then it follows that \(dx = 8 \cos\theta\,d\theta\). Furthermore, \[\begin{aligned}\sqrt{64-x^2} &= \sqrt{64-(8\sin \theta)^2}\\ &= \sqrt{64-64\sin^2\theta}\\ &= \sqrt{64(1-\sin^2\theta)} \\ &= \sqrt{64\cos^2\theta}\\ &= 8\cos\theta\end{aligned}\] Therefore, \(\begin{aligned}\int \sqrt{64-x^2}\,dx &= \int 8\cos\theta\cdot8\cos\theta\,d\theta\\ & = 64\int\cos^2\theta\,d\theta \\ &= 32\int 1+\cos(2\theta)\,d\theta \\ &= 32\theta + 16\sin(2\theta)+C\\ &=32\theta+32\sin\theta\cos\theta +C\end{aligned}\) At this point, you need to convert everything back to \(x\). You now need to use the following triangle to help you:  Do you think you can wrap things up from here? :-)

OpenStudy (anonymous):

\[\int\limits \sqrt{64(1-\sin ^{2}\theta)}dx\]

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

gimme a min

OpenStudy (anonymous):

rings a bell now i bet right?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

actually the original problem was \[\int\limits_{-8}^{8}\sqrt{64-x ^{2}}\]

OpenStudy (anonymous):

so i think i'll just change the limits accordingly

OpenStudy (anonymous):

Oh wow. Well, if that's the case, then this is super trivial if you recall what kind of shape \(\sqrt{64-x^2}\) is; then you can use the corresponding geometry formula to find the value of the integral. XD

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

wow i am dumb

OpenStudy (anonymous):

semicircle with radius 8

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

oh my all that for the areas of half a circle you can still do it the other way, but i am going to guess that this comes in the book long before the stupefyingly dull techniques of integration

OpenStudy (anonymous):

oh wait it's the average value so the answer is 2pi

OpenStudy (anonymous):

at least it was good practice

OpenStudy (anonymous):

that's all i can say

OpenStudy (anonymous):

Yes, good practice for what (not) to look forward to in calculus. XD

OpenStudy (anonymous):

:) Thanks a ton, both of you guys! I wish I could give 2 best responses.